OpenStudy (anonymous):

If "a" vector = 2 i+3j+k a.b=1 axb=j-k Find b vector..

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

pehchan kaun? :D

hartnn (hartnn):

someone who knows hindi :P

OpenStudy (anonymous):

okay ques btado

hartnn (hartnn):

take b as xi+yj+zk

OpenStudy (anonymous):

if we take cross with a?

OpenStudy (anonymous):

if we assume xi+yj+zk then what?

hartnn (hartnn):

did u take cross with a ?

hartnn (hartnn):

then dot product 1 gives 2x+3y+z=1

OpenStudy (anonymous):

(xi+yj+zk)(2i+3j+k)=1 2x+3y+z=1 okay then?

hartnn (hartnn):

know how to find aXb ?

OpenStudy (anonymous):

determinant?  i(3z-y)-j(2z-x)+k(2y-3x)=j-k

OpenStudy (anonymous):

lamba nahi ho raha? they have taken cross with a in the question axb=j-k mein then we get (a.b)a thing and we are given its value

hartnn (hartnn):

naah, thats what you need to do, solve the determinant and equate it to j-k

OpenStudy (anonymous):

then how to solve further?

hartnn (hartnn):

what you got for the determinant.? then you compare the values with i,j,k on both sides , value with i on left = 0 value with j = 1 value with k =-1

OpenStudy (anonymous):

i(3z-y)-j(2z-x)+k(2y-3x)=j-k 2z-x=1...(i)..........x=2z-1 3z-y=0.......y=3z...(ii) 2y-3x=-1...(iii) 6z-3x=-1 6z-3(2z-1)=-1 6z-6z+3=-1 z=0 :o

OpenStudy (anonymous):

getting weird values ~_~

hartnn (hartnn):

why can't 'z' be 0 ? you have the answer already ? i didn't check the calculation though...

OpenStudy (anonymous):

i have the answer and z not equal to 0

OpenStudy (anonymous):

@hartnn we will compare j to -1 i guess

OpenStudy (anonymous):

b is a vector @Loser66

hartnn (hartnn):

aXb = j-k so, compare j to 1 only ....not -1

OpenStudy (anonymous):

oh you're giving xyz?

OpenStudy (anonymous):

\[\Large b=\frac{1}{14}(6i+j-k)\] x should be 3/7 y should be 1/14 k should be -1/14 according to this

hartnn (hartnn):

2z-x=1...(i)..........x=2z-1 3z-y=0.......y=3z...(ii) 2y-3x=-1...(iii) are actually ONLY 2 independent equations you cannot solve them unless you take 2x+3y+z=1 into consideration then you have 3 independent equation, 3 unknowns

hartnn (hartnn):

2x+3y+z=1 y=3z x=2z-1 4z-2 +9z +z =1 14z = 3 z=3/14 samja ?

OpenStudy (anonymous):

but z should be -1/14

hartnn (hartnn):

abee yaar.... -j (2z-x) ha, right = j so, x-2z = 1 hoga na

hartnn (hartnn):

2x+3y+z=1 y=3z x=2z+1 4z+2 +9z +z =1 14z = -1 z=-1/14

hartnn (hartnn):

y= 3z = -3/14

OpenStudy (anonymous):

wahi to bola tha :/ -j se compare karenge

OpenStudy (anonymous):

6z-3x=-1<<<<<<NO aisa kyun par

hartnn (hartnn):

thats correct, i saw y=3z afterwards....

OpenStudy (anonymous):

okay then evt is okay but if we take axb=j-k iska cross with a then evt reduces to simple smhw

OpenStudy (anonymous):

a x a xb (a.b)a-(a.a)b a-14b now just compute axb and equate to a-14b

OpenStudy (anonymous):

(a.b)a = a..hmm

OpenStudy (anonymous):

a.b to 1 hai

hartnn (hartnn):

so, a=14b = aX (j-k)

hartnn (hartnn):

a-14b**

OpenStudy (anonymous):

but (a.b)a ko kaise khola

hartnn (hartnn):

a.a bhi 14 hai, 4+9+1=14

OpenStudy (anonymous):

yes a.a=14 and a.b=1

OpenStudy (anonymous):

(a.b)a-(a.a)b a - 14b

OpenStudy (anonymous):

latter one is okay but former me confusion hai

hartnn (hartnn):

whats the confusion xactly ?

OpenStudy (anonymous):

how did you open (a.b) a like a.a+ab smt?

hartnn (hartnn):

no! a.b =1 is given :P (a.b)a = 1a =a

OpenStudy (anonymous):

omg :/

OpenStudy (anonymous):

okay another thing when i say take cross with "a" aren't we doing in on the whole equation? like we divide the whole equation by a no. etc we wont take (j-k)xa or smt?

hartnn (hartnn):

you did aXaXb right ? so you crossed with a in the beginning, so it will be aX (j-k) ............NOT (j-k)Xa

hartnn (hartnn):

see if you get aX(j-k) as -4i +2j +2k i did it on paper

OpenStudy (anonymous):

yes wahi hai

hartnn (hartnn):

then a-14b = that find a-14b then compare i, j , k

OpenStudy (anonymous):

ha hogaya thanks ^.^

hartnn (hartnn):

welcome ^_^